What is the rate of consumption of O2 in a combustion reaction carried out in a 12.00 L flask if 7.42 mmol of oxygen is consumed in 13.9 s? Rate = 4.45e-5 Incorrect: Your answer is incorrect. M/s
Answer #1

Change in concentration of O2 = moles/volume

= 7.42×10-3/12 = 6.18×10-4 M

Rate of consumption = change in concentration/time taken

= 6.18×10-4/13.9

= 4.45×10-5 M/s